# Local EGOP for Continuous Index Learning

### v2

@@ 6.3 Noisy Manifold
Let $g=\nabla f(0)$, $H=\nabla^{2}f(0)$, $p=\pi(0)$, $\mathcal{T}:=\mathcal{T}_{p},\mathcal{N}:=\mathcal{N}_{p}$, and $T,R$ be as in Lemma 4. There exist non-negative constants $C_{\mathcal{N}}$ and $C_{\mathcal{T}\mathcal{N}}$ such that, for all $A\succeq 0$,
@@
We will leverage this weak orthogonal dependence to show that the distribution $\mu$ elongates along the space $\mathcal{N}_{x^{*}}$ orthogonal to the manifold. However, while the function is constant on this fiber, its gradient is not. Indeed, we are only guaranteed that the gradient remains tangent. Due to the effects of curvature, $\nabla f(p+v)$ may shrink or grow along different coordinate directions. We characterize this precisely as follows.
@@
Let $p\in\mathcal{M}$, $\eta\in\mathcal{N}_{p}$, $\|\eta\|<\tau$, $\tau$ the reach of $\mathcal{M}$. Then,
@@
In particular, for any such $\eta$, there exists a subspace $\operatorname{Null}=\operatorname{Null}_{\eta}$ of dimension at least D2dD-2d such that for any $v,z\in\operatorname{Null}$, $\nabla f(p+\eta+v)=\nabla f(p+\eta)$, $\nabla^{2}f(p+\eta+v)z=0$.
@@
Let $x^{*}=p+\eta$, $v\in\operatorname{Null}_{\eta}$. Define $\mu_{v}$ as the measure $\mu=N(x^{*},\Sigma)$ conditioned on $\pi_{\operatorname{Null}}[X-x^{*}]=v$, i.e. the draws where the $\operatorname{Null}$ component is fixed at $v$. Let $\Sigma_{v}$ be the covariance of this distribution. Then,
@@
where $\operatorname{Null}\subseteq\operatorname{nullspace}(H_{v})$, $\sup_{v\in\operatorname{Null}}\|R_{v}(A)\|\leq C_{\operatorname{Null}}\|A\|^{2},\ T_{v}(A)\leq C_{\operatorname{Null}}\|A\|$ for a uniform constant $C_{\operatorname{Null}}$, and $H_{v},T_{v},R_{v}$ are continuous in $v$, Lipschitz in $\|A\|$.
@@
Assume that $D\geq 2d$ and $\operatorname{rank}(H)=2d$. Let $\Sigma_{i}$ satisfy the recurrence in equation 7, $\sqrt{t_{i}}=\alpha r^{i}$, $0<r<1,\ \beta>0$. Then, for $\alpha,\zeta$ sufficiently small, $g^{T}\Sigma_{i}g=\Theta(t_{i})$, for $v\in\operatorname{Null}$, $v^{T}\Sigma_{i}v=\Theta(1)$, and for $\eta\in\operatorname{Null}^{\perp}\cap g^{\perp}$, $\eta^{T}\Sigma\eta=\theta(\sqrt{t_{i}})$. It follows that, selecting $i_{n}=\frac{4\log n}{(2d+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{\Sigma_{i_{n}}}(f)-f)^{2}\,d\mu_{i_{n}}\right]=O\left(n^{-\frac{4}{2d+5}}\right).$
Rather than fully denoising the manifold, our analysis reveals that the manifold tangent, as well as the normal vectors that do not fall in $\operatorname{Null}$, are subject to the same second-order anisotropy as observed in the full rank setting. These compose precisely those directions in which the gradient varies at first order, leaving only the inactive subspace where the linearization is preserved. Practically, the resulting rate is comparable to that of a standard estimator applied to a completely denoised manifold, as this method yields an effective dimension of $d+1/2$.
Example 3 (Spheres).
$H=\begin{bmatrix}T&b\\ b^{T}&0\end{bmatrix}$
@@
Figure 5: Eigenvalue decay of $\Sigma_{i}$ with Local EGOP Learning applied to data satisfying the noisy manifold hypothesis about $S^{2}$. The orthogonal exhibits light decay, approaching second order anisotropy asymptotically.
Example 5 (Multi-Index Learning).
When the underlying manifold $\mathcal{M}$ is flat, the supervised noisy manifold hypothesis reduces to a multi-index setting. The key distinction is that displacement away from the manifold not only preserves the function value, but also all higher order function derivatives. In particular, the Hessian has rank no more than $d$, the active index dimension, and the Hessian column space is shared across the distribution.
Thus, when performing expansions such as Corollary 1, the orthogonal has no contribution to the EGOP, leading to implicit dimension reduction. In particular, this yields a learning rate of $(d+1)/2$.
@@ 7 Experiments
We provide the following simulations to empirically justify our theoretical results. First, expanding on the noisy manifold setting, we show how the regression error rates of Local EGOP Learning remain invariant to the injection of high-dimensional noise. We demonstrate that two-layer neural networks are not able to efficiently learn in the noisy manifold setting, with a sharp decrease in performance compared to Local EGOP Learning. Next, we compare the performance of our method to that of a deep neural network on toy data. In particular, we show how the feature embeddings produced by transformers trained on a simple example are qualitatively similar to the localizations generated by our procedure. Finally, we apply this procedure to estimate backbone angles in Molecular Dynamics (MD) data, where we leverage the noisy manifold structured data to improve prediction quality.
We relegate details for the evaluation and fitting of each method to Appendix A.
Figure 6: (Left) Comparison of Local EGOP Learning to the performance of a two-layer neural network architectures trained on helical data in ambient dimension $D=5$. (Right) Local EGOP Learning trained on helical data in various ambient dimensions $D$.
@@ 7.1 Learning Rate
7.2 Two-layer Neural Network
We assess a variety of two-layer neural network architectures, with implementation and training explained in detail in Appendix A.
Helical data provides a rich experimental setting to compare Local EGOP Learning to two-layer neural networks. It is not only continuous single-index, it is multi-index 2 for $D$ even, and 3 for $D$ odd. Further, by construction it is well-approximated by a low intrinsic dimensional manifold, a typical setting for accelerated learning rates (Kiani et al. (2024); Liu et al. (2021), etc.).
Unsurprisingly, as demonstrated in Figure 6, the two-layer neural networks readily adapt to the multi-index data structure. For $D=5$, the overparameterized ReLU network achieves comparable error to a multi-index architecture pre-specified to learn a 3-dimensional subspace. However, Local EGOP Learning vastly reduced the error for moderate to large sample sizes. This suggests that no architecture was able to fully adapt to the intrinsic continuous single-index structure.
@@ 7.4 Predicting the backbone angles in Molecular Dynamics (MD) data
Our final example comes from the analysis of molecular geometries. The raw data consist of $X,Y,Z$ coordinates for each of the $N_{a}$ atoms of a molecule, which, due to interatomic interactions, lie near a low-dimensional manifold (Das et al., 2006). While the governing equations of the simulated dynamics are unknown, for small organic molecules, certain backbone angles (Das et al., 2006) have been observed to vary along the aforementioned low-dimensional manifold. Specifically, for the malonaldehyde molecule, the two backbone angles denoted $\tau_{1,2}$ are shown in Figure 7(b). We used a subsample of molecular configurations of size $n=10^{4}$ from the MD simulation data of Chmiela et al. (2017) as input data. The configuration data, pre-processed as in Koelle et al. (2022), consists of $D=50$ dimensional vectors and lie near a 2-dimensional surface with a torus topology (see Figure 7(b)). On a hold-out set of 500 test points, Local EGOP Learning yields an MSE of 0.00043, compared to 0.012 for Gaussian kernel Nadaraya-Watson with cross-validated bandwidth selection.
@@ 8 Discussion and Future Work
The primary contribution of this work is to establish a rigorous basis for the study of recursive kernel machines that localize about points of interest. We leverage this framework to verify that Local EGOP Learning achieves desirable learning rates in model settings. Of particular importance is our introduction of continuous index learning, formalized by the supervised noisy manifold hypothesis, under which anisotropy about the orthogonal is achieved under appropriate non-degeneracy assumptions. Our algorithm is most similar to the heuristic Takeda et al. (2007), hence we believe that the present framework can be extended to analyze this algorithm, too. Local EGOP takes its name from Radhakrishnan et al. (2022), with the key distinction being that the AGOP is averaged over the localized region $\mu_{i}$ rather than the entire data distribution $P_{n}$.
We demonstrate the efficacy of this algorithm, qualitatively when compared to a transformer architecture, and quantitatively on high dimensional continuous single-index data. The theoretical framework developed in this work opens many avenues for future study. In addition to the open problems this investigation has raised, our approach is conducive to a variety of strategies that may further improve the stated results. We mention several of the most relevant directions.
For one, we can consider optimizing the objective
$\operatorname{tr}[\mathcal{L}(\mu_{t})\Sigma(\mu_{t})]+\frac{1}{n\sqrt{\det\Sigma(\mu_{t})}}$
in a variety of ways. Working with a second order relaxation, one may seek to optimize the covariance gradually, such as by a gradient-based method. Alternatively, motivated by Proposition 1, one can show that this ratio is minimized by a distribution $\mu_{t}$ such that $\Sigma(\mu_{t})=t\mathcal{L}(\mu_{t})^{-1}/D$. Differentiating, this yields a differential equation
$\partial_{t}\mathcal{L}(\mu_{t})\Sigma(\mu_{t})+\mathcal{L}(\mu_{t})\partial_{t}\Sigma(\mu_{t})=I/D$
@@
which may be amenable to analysis, particularly in the Gaussian setting. However, it may also be fruitful to drop both this assumption and the covariance relaxation. One promising direction is to view $W(\mu_{t})$ and $\Sigma(\mu_{t})$ as functionals, and to then optimize the ratio using the Wasserstein gradient flow (Ambrosio et al., 2006) or other methods that operate at the distribution level. In particular, this would allow for non-ellipsoidal localizations, potentially allowing for improved learning rates in settings where the level-set is curved. A fundamental hurdle for any of these approaches is that, generally, the EGOP-form can only upper bound localization-induced bias, and the resulting bound is not always sharp.
Many basic questions have been left for future research. As discussed in Appendix C, in the context of Local EGOP Learning, the problem of EGOP estimation is left open. Algorithm 1 is slow, requiring many local linear regressions at each iteration. We presented this simple algorithm for clarity, and many practical improvements can be made to improve its performance, particularly when multiple data points are queried for prediction. Furthermore, simple modifications to the algorithm would yield gains in statistical efficiency. Higher-order local polynomial regression can be performed for function estimation. This would improve theoretical guarantees by accounting for more general smoothness patterns in the function of interest, but it may also inform more complex localization procedures. Gaussian localizations are quadratic in nature, and so a more generic formulation may yield increased adaptivity.
Importantly, Local EGOP Learning is based firmly on a MSE formulation of regression, restricting its utility to $\ell_{2}$ loss. It is unlikely that this method can be applied directly to problems such as classification or image generation. This leaves open how an appropriate adaptation should be formulated in these settings. Additionally, a formal analysis comparing procedures such as Local EGOP Learning to modern neural network architectures is merited. In particular, while we do not expect an exact replication of behavior, the Local EGOP ansatz, which asserts that learned features ought to spread locally inversely proportionate to the EGOP, may inform further study of these learning algorithms, or lead to improvements in the design of model architectures.
@@ Appendix A Experiment Details
(b) Coordinates of Malonaldehyde embedded into 2 dimensions via Diffusion Maps [Coifman and Lafon, 2006] and colored by the backbone angles $\tau_{1}$.
@@ A.1 General Set-up
where $c_{\alpha}\sim\mathrm{Uniform}[-3,3]$. The resulting function $f$ is then evaluated at the unperturbed datapoints before the features $X$ were contaminated with noise orthogonal to the underlying manifold $\mathcal{M}$. The coefficients $c_{\alpha}$ are randomly sampled across each replicate of the experiments. We generically apply the below hyperparameters to the Local EGOP Learning algorithm in each example.
@@
  • Bandwidth decay $t_{i}=(1+i)^{-1.2}$,
  • Initialization scale $\alpha=0.2$
  • Exclusion radius $\varepsilon=1.6$
@@ A.2 Two-layer Neural Network
For moderate-scale neural models, we conduct a grid search over hyperparameters. Widths are selected from $128,256,512,1024,2048,4096,8192$, learning rates are selected from $10^{-3},3\times 10^{-3}$, weight decay from $0,10^{-5},10^{-4}$, and batch size from $16,32,64$. The default Adam momentum parameters are used. Early stopping is applied with patience $200$ and improvement tolerance $10^{-4}$ monitored on a validation split. In each experiment, only the optimal test error of these methods is aggregated into the “Small Tuned NN” plot.
We train overparameterized two-layer models with width $100,000$, Adam with learning rate $0.003$, batch size $64$, and fixed weight decay $10^{-4}$, applying early stopping with patience $300$ and improvement tolerance $10^{-4}$.
We also consider SIL models $f(x)=g(\beta^{\top}x)$, $g$ a two-layer NN, with width 256 and GELU activations, $\beta$ is normalized at each iteration.
@@ A.3 Feature Learning
To qualitatively compare the local weighting behavior of EGOP with that of a learned neural embedding, we constructed a unified visualization pipeline producing animated opacity maps over a shared dataset. Both models were applied to the same two-dimensional manifold samples, with visualization frames showing how point-wise influence weights evolve during training or iteration.
Setup.
We generated n=50,000n=50{,}000 samples on a spherical cap in $\mathbb{R}^{2}$, restricted to an angular width of $25^{\circ}$. Each point $x_{i}$ was scaled radially by a random factor $r_{i}\sim\mathrm{Uniform}[r,R]$, with $r=0.6$ and $R=1.8$, and colored by the output of a random cubic polynomial $f(x)$. Additive Gaussian noise with standard deviation $0.05$ was applied to the responses. The cap’s pole $x_{0}=(1,0)^{\top}$ was used as a fixed reference point for visualization of weight evolution.
We trained an FTTransformer regression model [Gorishniy et al., 2021] on this dataset using the rtdl and zero frameworks. The network was trained for 40 epochs with a batch size of 1024 and learning rate $2\times 10^{-3}$. After every 50 training steps, we computed an embedding-based influence profile analogous to EGOP weights:
$w_{i}=\frac{\exp(-\|z_{i}-z_{0}\|^{2}/\tau)}{\sum_{j}\exp(-\|z_{j}-z_{0}\|^{2}/\tau)},$
where $z_{i}$ denotes the learned feature representation of $x_{i}$, $z_{0}$ is the embedding of the reference point, and $\tau=8.0$ is a temperature parameter controlling decay. These weights were rendered with the same opacity scheme as EGOP to enable direct visual comparison.

### v3

@@ 4.4.3 Noisy Manifold
Let $g=\nabla f(x^{*})$, $H=\nabla^{2}f(x^{*})$, $p=\pi(x^{*})$, $\mathcal{T}:=\mathcal{T}_{p},\mathcal{N}:=\mathcal{N}_{p}$, and $T,R$ be as in Lemma 3. There exist non-negative constants $C_{\mathcal{N}}$ and $C_{\mathcal{T}\mathcal{N}}$ such that, for all $A\succeq 0$,
@@
We will leverage this weak orthogonal dependence to show that $\mu_{t}$ elongates along the space $\mathcal{N}_{x^{*}}$ orthogonal to the manifold. However, while the function is constant on this fiber, its gradient is not. Indeed, we are only guaranteed that the gradient remains tangent. Due to the effects of curvature, $\nabla f(p+v)$ may shrink or grow along different coordinate directions. We characterize this precisely as follows.
@@
Let $p\in\mathcal{M}$, $\eta\in\mathcal{N}_{p}$, and $\|\eta\|<\tau$, where $\tau$ denotes the reach of $\mathcal{M}$. For $S_{p}$ the shape operator,
@@
For any such $\eta$, there exists a subspace $\operatorname{Null}=\operatorname{Null}_{p+\eta}$ of dimension at least D2dD-2d such that, for any $v,z\in\operatorname{Null}$, $\nabla f(p+\eta+v)=\nabla f(p+\eta)$, $\nabla^{2}f(p+\eta+v)z=0$.
@@
Let $x^{*}=p+\eta$, $v\in\operatorname{Null}_{p+\eta}$. Define $\mu_{v}$ as the measure $\mu=N(x^{*},\Sigma)$ conditioned on $\pi_{\operatorname{Null}}[X-x^{*}]=v$, i.e. the draws where the $\operatorname{Null}$ component is fixed at $v$. Let $\Sigma_{v}$ be the covariance of this distribution. Then,
@@
where $\operatorname{Null}\subseteq\operatorname{nullspace}(H_{v})$, $\sup_{v\in\operatorname{Null}}\|R_{v}(A)\|\leq C_{\operatorname{Null}}\|A\|^{2},\ T_{v}(A)\leq C_{\operatorname{Null}}\|A\|$ for a uniform constant $C_{\operatorname{Null}}$, and $H_{v},T_{v},R_{v}$ are continuous in $v$ and Lipschitz in $\|A\|$.
@@
Assume that $D\geq 2d$ and $\operatorname{rank}(H)=2d$. Let $\Sigma_{t}$ satisfy the recurrence in equation 2, $\sqrt{h_{t}}=\alpha r^{t}$, $0<r<1,\ \beta>0$. Then, for $\alpha,\zeta$ sufficiently small, $g^{T}\Sigma_{t}g=\Theta(h_{t})$. If $v\in\operatorname{Null}$, then $v^{T}\Sigma_{t}v=\Theta(1)$, and if $\eta\in\operatorname{Null}^{\perp}\cap g^{\perp}$, $\eta^{T}\Sigma\eta=\Theta(\sqrt{h_{t}})$. It follows that, selecting $t_{n}=\frac{4\log n}{(2d+5)\log(1/r)}$,
$\mathbb{E}\left[\int(\hat{P}_{M_{t_{n}}}(f)-f)^{2}\,d\mu_{t_{n}}\right]=O\left(n^{-\frac{4}{2d+5}}\right).$
Rather than fully denoising the manifold, our analysis reveals that the manifold tangent, as well as the normal vectors that do not fall in $\operatorname{Null}$, are subject to the same second-order anisotropy as observed in the full rank setting. These compose precisely those directions in which the gradient varies at first order. The resulting rate is comparable to that of a standard estimator applied to a completely denoised manifold, yielding an effective dimension of $d+1/2$.
Example 3 (Multi-Index Learning).
When the underlying manifold $\mathcal{M}$ is flat, the supervised noisy manifold hypothesis reduces to a multi-index setting. The key distinction is that displacement away from the manifold not only preserves the function value, but also all higher order function derivatives. In particular, the Hessian has rank no more than $d$. When performing expansions such as Corollary 1, the orthogonal has no contribution to the EGOP, yielding a learning rate of $(d+1)/2$.
@@ 5 Experiments
Details for data generation, evaluation, and training of each method can be found in Appendix A.
Figure 5: (Left) Comparison of Local EGOP Learning to the performance of two-layer neural network architectures trained on helical data in ambient dimension $D=5$. (Right) Local EGOP Learning trained on helical data in various ambient dimensions $D$.
5.1 Two-layer Neural Network
We assess a variety of two-layer neural network architectures, with details in Appendix A. The covariate distribution is drawn from helical data, which is continuous single-index, being drawn from a noise contaminated curve, as well as multi-index 2 for $D$ even, and 3 for $D$ odd. Further, by construction it is well-approximated by a low intrinsic dimensional manifold, a typical setting for accelerated learning rates (Kiani et al. (2024); Liu et al. (2021), etc.).
As demonstrated in Figure 5, the two-layer neural networks readily adapt to the multi-index data structure. For $D=5$, the overparameterized ReLU network achieves comparable error to a multi-index architecture pre-specified to learn a 3-dimensional subspace. However, Local EGOP Learning vastly reduces the error for moderate to large sample sizes. This suggests that no architecture was able to fully adapt to the intrinsic continuous single-index structure.
@@ 5.4 Predicting the backbone angles in Molecular Dynamics (MD) data
Our final example comes from the analysis of molecular geometries. The raw data consist of $X,Y,Z$ coordinates for each of the $N_{a}$ atoms of a molecule, which, due to interatomic interactions, lie near a low-dimensional manifold (Das et al., 2006). While the governing equations of the simulated dynamics are unknown, for small organic molecules, certain backbone angles (Das et al., 2006) have been observed to vary along the aforementioned low-dimensional manifold. Specifically, for the malonaldehyde molecule, the two backbone angles denoted $\tau_{1,2}$ are shown in Figure 6(b). We used a subsample of molecular configurations of size $n=10^{4}$ from the MD simulation data of Chmiela et al. (2017) as input data. The configuration data, pre-processed as in Koelle et al. (2022), consist of $D=50$ dimensional vectors and lie near a 2-dimensional surface with a torus topology (see Figure 6(b)). On a hold-out set of 500 test points, Local EGOP Learning yields an MSE of 0.00043, compared to 0.012 for Gaussian kernel Nadaraya-Watson with cross-validated bandwidth selection.
@@ 6 Discussion and Future Work
The problem of learning an EGOP (Trivedi et al., 2014; Radhakrishnan et al., 2025), its local counterpart (Takeda et al., 2007), or gradients (Mukherjee et al., 2010; Wu et al., 2010), has been a topic of frequent theoretical interest, with applications to image processing (Takeda et al., 2007, 2008), machine learning (Radhakrishnan et al., 2022; Beaglehole et al., 2023), and dimension reduction (Yuan et al., 2023). Our research establishes a new functional framework to understand these methods, directly relating the EGOP to the MSE learning objective, and recursive kernel learning to a greedy variance reduction strategy. The resulting algorithm yields not only new efficiency guarantees for point estimation, but also tractable access to a local EGOP metrization.
In this paper, we introduce continuous index learning and the closely related supervised noisy manifold hypothesis. We demonstrate that Local EGOP Learning is able to achieve an intrinsic dimensional learning rate for such data. For generic manifolds, however, it does not completely denoise the covariate distribution. Our theory shows that a small subspace of the orthogonal will contract at the same rate as the tangent, despite the invariance of $f$. This counterintuitive result reflects the fact that our EGOP optimization scheme is derived via an upper-bound on the bias, and is thus not always sharp. The refinement of this methodology to close this gap is thus of both practical and theoretical importance.
The Local EGOP Learning algorithm is adaptive, in that it outperforms standard kernel regression algorithms when the noisy supervised manifold hypothesis fails to hold (Theorem 2), and yields an intrinsic dimensional rate when $f$ is structured. In neither case is knowledge of the regularity of $f$ or the latent manifold required to enhance estimation quality. Importantly, this means that users may use the algorithm without a priori knowledge of problem dimensionality, noise level, or whether the supervised manifold hypothesis holds.
Algorithm 1 was presented for clarity, not efficiency; there are many possible computational improvements, including partitioning schemes as in Beaglehole et al. (2025). Local EGOP Learning is based firmly on a MSE formulation of regression, restricting its utility to $\ell_{2}$ loss. This method requires suitable modification to be applied directly to problems such as classification or image generation. This leaves open how an appropriate adaptation should be formulated in these settings.
@@
Our methodology is primarily framed through a second order, Gaussian relaxation. One promising direction is to drop this restriction, and view $W(\mu_{t})$ and $\Sigma(\mu_{t})$ as functionals, and to then optimize the MSE using a Wasserstein gradient flow (Ambrosio et al., 2006) or other methods that operate at the distribution level. In particular, this would allow for non-ellipsoidal localizations, potentially allowing for improved learning rates in settings where the level-set is curved. A fundamental hurdle for any of these approaches is that, generally, the EGOP form can only upper bound localization-induced bias, and the resulting bound is not always sharp.
@@ Appendix A Experiment Details
(b) Coordinates of Malonaldehyde embedded into 2 dimensions via Diffusion Maps (Coifman and Lafon, 2006) and colored by the backbone angles $\tau_{1}$.
Figure 6: Visualizations of the experimental datasets.
@@ A.1 General Set-up
The following simulations to empirically justify our theoretical results. First, expanding on the noisy manifold setting, we show how the regression error rates of Local EGOP Learning remain invariant to the injection of high-dimensional noise. We demonstrate that two-layer neural networks are not able to efficiently learn in the noisy manifold setting, with a sharp decrease in performance compared to Local EGOP Learning. We next compare the performance of our method to that of a deep neural network on toy data. In particular, we show how the feature embeddings produced by transformers trained on a simple example are qualitatively similar to the localizations generated by our procedure. Finally, we apply this procedure to estimate backbone angles in Molecular Dynamics (MD) data, where we leverage the noisy manifold structured data to improve prediction quality.
@@
where $c_{\alpha}\sim\mathrm{Uniform}[-3,3]$. The resulting function $f$ is then evaluated at the unperturbed datapoints before the features $X$ were contaminated with noise orthogonal to the underlying manifold $\mathcal{M}$. The coefficients $c_{\alpha}$ are randomly sampled across each replicate of the experiments. In our plots, confidence bands are displayed with a width of $\pm 1$ standard error. We generically apply the below hyperparameters to the Local EGOP Learning algorithm in each example.
@@
  • Bandwidth decay $h_{t}=(1+t)^{-1.2}$,
  • Initialization scale $\alpha=0.2$,
  • Exclusion radius $\varepsilon=1.6$.
@@ A.2 Two-layer Neural Network
For moderate-scale neural models, we conduct a grid search over hyperparameters. Widths are selected from $\{128,256,512,1024,2048,4096,8192\}$, learning rates are selected from $\{10^{-3},3\times 10^{-3}\}$, weight decay from $\{0,10^{-5},10^{-4}\}$, and batch size from $\{16,32,64\}$. The default Adam momentum parameters are used. Early stopping is applied with patience $200$ and improvement tolerance $10^{-4}$ monitored on a validation split. In each experiment, only the optimal test error of these methods is aggregated into the “Small Tuned NN” plot.
We train overparameterized two-layer models with width 100,000100{,}000, Adam with learning rate $0.003$, batch size $64$, and fixed weight decay $10^{-4}$, applying early stopping with patience $300$ and improvement tolerance $10^{-4}$.
We also consider multi-index models $f(x)=g(Ux)$, $g$ a two-layer NN, with width 256 and GELU activations, $U$ a rank three projection matrix updated at each iteration.
@@ A.3 Feature Learning
We generated n=50,000n=50{,}000 samples on a spherical cap in $\mathbb{R}^{2}$, restricted to an angular width of $25^{\circ}$. Each point $x_{i}$, we evaluate $f(x_{i})$ for $f$ a cubic polynomial with $\mathrm{Uniform}[-1,1]$ coefficients. The features $x_{i}$ are then scaled radially by a random factor $r_{i}\sim\mathrm{Uniform}[r,R]$, with $r=0.6$ and $R=1.8$. Additive Gaussian noise with standard deviation $0.05$ was applied to the labels. We set $x^{*}=(1,0)^{T}$.
We trained an FTTransformer regression model (Gorishniy et al., 2021) on this dataset using the rtdl and zero frameworks. The network was trained for 40 epochs with a batch size of 1024 and learning rate $2\times 10^{-3}$. After every 50 training steps, we computed weights:
$w_{i}=\frac{\exp(-\|z_{i}-z^{*}\|^{2}/8)}{\sum_{j}\exp(-\|z_{j}-z^{*}\|^{2}/8)},$
where $z_{i}$ denotes the learned feature representation of $x_{i}$.
A.4 Unsupervised Embedding
As noted in works such as (Coifman and Lafon, 2006; El Karoui, 2010; Landa and Cheng, 2023; Kokot et al., 2025), unsupervised methods such as diffusion maps are robust to noise. The resulting low-frequency eigenfunctions closely match those of the underlying manifold. However, as demonstrated in Figure 7, they fail to capture this structure exactly, and thus cannot match the metrization produced by Local EGOP Learning and deep learning architectures in Figure 3(a).
Figure 7: Data sampled from an annulus, with red rings displaying uniform distances in diffusion map embedding distance from the point $(1,0)^{T}$, for a 2-dimensional embedding.